OpenStudy (anonymous):

f(x) = 5x^2 x = 10 find the derivative algebraically

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

What do you mean by that? Something like this? \[ f'(10)=\lim_{\Delta x->0} \frac{f(10+\Delta x)-f(10)}{\Delta x}\\ \ \ \ \ \ \ \ =\lim_{\Delta x->0} \frac{5(10+\Delta x)^2-500}{\Delta x}\\ \ \ \ \ \ \ \ =\lim_{\Delta x->0} \frac{5(100+20\Delta x+{\Delta x}^2)-500}{\Delta x}\\ \ \ \ \ \ \ \ =\lim_{\Delta x->0} \frac{500+100\Delta x+5{\Delta x}^2-500}{\Delta x}\\ \ \ \ \ \ \ \ =\lim_{\Delta x->0} \frac{100\Delta x+5{\Delta x}^2}{\Delta x}\\ \ \ \ \ \ \ \ =\lim_{\Delta x->0} [100+5{\Delta x}]\\ \ \ \ \ \ \ \ =100 \]

jimthompson5910 (jim_thompson5910):

The derivative of f(x) using the limit definition is \[\Large \lim_{h\to0} \frac{f(x+h)-f(x)}{h} \]

jimthompson5910 (jim_thompson5910):

So we need to find f(x+h) to go forward

jimthompson5910 (jim_thompson5910):

f(x) = 5x^2 f(x+h) = 5(x+h)^2 ... replace all copies of 'x' with 'x+h' Then simplify/expand f(x+h) = 5(x+h)(x+h) f(x+h) = 5[x(x+h) + h(x+h)] f(x+h) = 5[x^2+xh + hx+h^2] f(x+h) = 5[x^2+ 2xh+h^2] f(x+h) = 5x^2+ 10xh+5h^2

OpenStudy (anonymous):

\[\lim x->0 5(x ^{2} +2hx + h ^{2}) -5h ^{2} / h \]